OpenStudy (anonymous):

1. On the grid below, solve the system of equations graphically for x and y. 4x-2y=10 y=-2x-1

jimthompson5910 (jim_thompson5910):

2x + y = 1 y = 1 - 2x -------------------- 4x + 2y = -1 4x + 2(1 - 2x) = -1 4x + 2 - 4x = -1 I'll let you finish up